OpenStudy (anonymous):

Find the value of y for a given value of x, if y varies directly with x. If y = 39 when x = –117, what is y when x = –132?

OpenStudy (anonymous):

-3y=x -3y=132 y=-44

OpenStudy (anonymous):

y=39 x=-117 y when x=-132 39/-117=k k=-0.3 y=-0.3*-132 y=39.6 thats what i got,

OpenStudy (anonymous):

but k don't equal -0.3.

OpenStudy (anonymous):

it equals -1/3

OpenStudy (anonymous):

oh i still not getttin -44 tho

OpenStudy (anonymous):

39/-117=k k=1/3 y=1/3*-132 y=-44

OpenStudy (anonymous):

put this in a calculator (1/3) x -132 = and you will get -44